A large and growing retail chain is opening a new warehousing operation that will operate 8

hours per day. They expect 256 trucks to drop off and then pick up products each weekday.

Trucks enter each loading dock one

at a time and it takes 50 minutes to unload and then refill a

truck before it can leave. Each truck carries an average of 24 pallets of merchandise at any one

time.

Truck teams of two people are responsible for moving merchandise on and off of the trucks

and

tracking inventory received and sent. Each of these teams is assigned to a loading dock. For

every three truck teams, there is one extra employee that will cover breaks and help coordinate

work.Lift trucks are used to pick up merchandise from the loading docks and place it in its proper

storage area of the warehouse. On their return, these lifts bring other merchandise from the

warehouse and deliver it to the loading docks to later be placed on trucks for delivery. It takes an

average of 2.5 minutes for a lift truck to transport merchandise between the loading dock and the

storage areas (one way). Lifts are not assigned to specific loading docks and each can carry only

one pallet at a time.

Using the Naïve approach, calculate the minimum requirements for:

a)

The number of loading docks that need to be built (Facility requirements)

b)

The number of Lift trucks needed (Equipment requirements)

c)

The number of employees working as part of the truck teams (Labour requirements)

d)

How many more trucks could be serviced (if any) without having to build more loading

docks?
Answer #1

Number of trucks available - 256 trucks/ day

Number of hours per day - 8

Average time taken to refill and unload a truck - 50 minutes/ truck

Average number of pallets carried by each truck - 24 pallets

1) Number of docks required :

Number of trucks serviced in an 8-hour shift per dock

= 8 hours * 60 minutes / (50 minutes per truck + 2.5 minutes to transport)

= 480 / 52.5 = 9 trucks

For 256 trucks, number of docks required = 256 docks / 9 = 28.4 docks = 28 docks

2) Number of lift trucks needed :

Each dock has one lift truck. Hence, number of docks = Number of lift trucks

Number of lift trucks needed = 28

3) Number of employees working on the truck team :

As per the case, number of people working on one truck = 2

One person additional for every 3 trucks.

Total number of employees = 2 * 9 + 3 (one employee per 3 trucks)

= 21 employees